# A Pre-trained Sequential Recommendation Framework: Popularity Dynamics for Zero-shot Transfer

### 03 Jan 2024 v1

@@ 5.1. Datasets and Preprocessing
5.2. Baselines and Experimental Setup
To assess the effectiveness of our proposed method, we compare against the following set of baselines. These baselines include classic general recommendation models (e.g., MostPop, BPR (Rendle et al., 2009), NCF (He et al., 2017), NGCF (Wang et al., 2019), LightGCN (He et al., 2020)) and state-of-the-art sequential recommendation models (e.g., Caser (Tang and Wang, 2018). MArank (Yu et al., 2019), SasRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019), TiSasRec (Li et al., 2020)). We provide more detailed descriptions of the baselines in Appendix B. To the best of our knowledge, we are the first to tackle the zero-shot sequential recommendation without any auxiliary information. Therefore, we do not include the existing pre-trained sequential recommenders (Hou et al., 2022; Ding et al., 2021) as they require textual descriptions of items, which we do not consider in the experiments.
Following previous works (Koren, 2008; He et al., 2017; Kang and McAuley, 2018; Sun et al., 2019), we adopt the leave-one-out evaluation method: for each user, we pair its test item with 100 unobserved items according to the user’s interaction history. Then we rank the test item for the user among the 101 total items. We use two standard evaluation metrics for top-$k$ recommendation: Recall@k (R@k) and Normalized Discounted Cumulative Gain@N (N@k). R@k measures the proportion of test items that appear in the top-$k$ recommendations, and N@k accounts for the ranking position of the test interaction. We report the average of R@k and N@k over all the test interactions.
We use publicly available Pytorch implementations for the baselines. For fair evaluation, we set dimension size $d$ to 50, max sequence length $L$ to $200$, and batch size to $128$ in all models. We use three layers (one base and two additional) in NCF and NGCF, and two attention layers for the sequential models (SasRec, BERT4Rec, TiSasRec, and ours). We use an Adam optimizer and tune the learning rate in the range $\{10^{-4},10^{-3},10^{-2}\}$ and set the weight decay to $10^{-5}$. We use the dropout regularization rate of $0.3$ for all models. We set $\gamma=0.5$ in Equation 1, whose reason we will discuss the reason in  § 5.5.1. We define the coarse and fine time period to be a month and a week respectively, and we fix the window size to be $m=12$ and $n=4$ for all datasets (§ 4.1.1). We train PrepRec for a maximum of 80 epochs. All experiments are conducted on a Tesla V100 using PyTorch. We repeat each experiment 5 times with different random seeds and report the average performance.
@@ 5.3. Regular Sequential Recommendation (RQ1)
We investigate the performance of PrepRec against the selected baselines. Note that the purpose of this paper is not to outperform the state-of-the-art sequential recommenders. Instead, we include the results here to investigate how well the popularity dynamics can capture users’ preferences in the sequential recommendation and can explain the state-of-the-art models. To no surprise, the selected sequential recommendation baselines are the best performant. However, PrepRec still achieves competitive performance- within 5% in R@10, - with the state-of-the-art sequential recommenders. On Epinions, PrepRec even outperforms all the sequential recommenders by 7.3%. We argue that this is particularly impressive given that PrepRec has significantly fewer model parameters (Table 9) and is capable of zero-shot transfer, with no side information.
@@
PrepRec learns item representations through popularity dynamics, which is conceptually different from learning representations specific to each item ID. Therefore, we propose a simple post-hoc interpolation to investigate if modeling popularity dynamics can boost the performance of state-of-the-art sequential recommenders. We interpolate the scores from PrepRec with the scores from BERT4Rec as follows: $\hat{y}_{intp}(v^{t}_{j}|\mathcal{S}_{u})=\alpha*\hat{y}_{O}(v^{t}_{j}|\mathcal{S}_{u})+(1-\alpha)*\hat{y}_{S}(v_{j}|\mathcal{S}_{u})$ where $\hat{y}_{O}(v^{t}_{j}|\mathcal{S}_{u})$ and $\hat{y}_{S}(v_{j}|\mathcal{S}_{u})$ are the scores from PrepRec (Equation 10) and BERT4Rec, respectively. We then proceed to rank with the final interpolated scores. We set $\alpha=0.5$ for all datasets. After interpolation, there is a significant performance boost by up to 39.7% in N@10. This suggests that our model effectively captures dynamic item properties that are missing with the static item embeddings of SOTA sequential models. Gains are largest in the medium and low-density datasets (Epinions, Amazon), indicating that our model is particularly complimentary to existing methods in sparse datasets where item embeddings are less informative.
5.4. Cross-domain Zero-shot Transfer(RQ2)
For the zero-shot transfer experiments, we train PrepRec on the source dataset. We follow the zero-shot inference setting introduced in § 4.3. We make predictions using Equation 12 and report the results in Table 4. We also include the results of the regular sequential recommendation for reference. In the zero-shot setting, after training on source datasets, in the target dataset, PrepRec shows minimal performance reduction in most cases (i.e., 2.3% maximum reduction) and even outperforms PrepRec and selected sequential recommendation baselines that are trained on the target datasets. In particular, we test all combinations of source and target pairs. We find the zero-shot transfer result from Tool to Office (i.e., Tool is the source dataset while office is the target dataset) outperforms all the state-of-the-art sequential recommenders trained on office, i.e., 0.548 vs 0.541 (BERT4Rec) in R@10. Similarly, when the target is Music and Epinions, using Movie as the source dataset attains nearly identical performance with the reference models. We conjecture that source datasets with similar density levels to the target dataset generally achieve excellent generalization performance, indicating that our model captures relevant characteristics across the space of datasets. Overall, these results show our model’s effectiveness in transferring across domains and applications, performing at or above par with SOTA models trained on the target dataset, without any training on interaction data or side information. We provide a more detailed analysis (case study) in Appendix F.
@@ 5.5. Ablation Study (RQ3)
5.5.1. Effect of discounting factor $\gamma_{a}$
We examine the effect of different preprocessing weights $\gamma$ used in popularity percentile calculation. In particular, $\gamma=1$ corresponds to the cumulative popularity, or in other words, at a given time period $t$, the overall number of interactions up to period $t$. On the other hand, $\gamma=0$ corresponds to the current popularity, or percentiles are calculated over interactions just in $t$, same as $b_{j}^{t}$ in Equation 1. When $\gamma=0.5$, it can be interpreted as interactions being exponentially weighted by time, with a half-life of 1 time period. We find that $\gamma=0.5$ generally outperforms the other two settings, with the largest gains of around 12% R@10 and 22% N@10 over cumul-pop in the dense Music dataset, and the largest gains over curr-pop in the sparser Office (4% N@10) and Epinions ($9\%$ R@10 and N@10) datasets. We suspect this is due to cumulative measures in denser datasets failing to capture recent trends due to the large historical presence, while current-only measures in sparser datasets convey too little or noisy information and lose the information of long-term trends. curr-pop shows decent performance on the Music dataset, suggesting that Music trends might be more cyclical and thus the current popularity is more informative.
@@ 5.5.2. Effect of Prediction Time
In PrepRec, we use the popularity trends up to some time as the context to construct universal item representations. To prevent information leakage, we can only use the popularity dynamics computed prior to the time of the interaction, as illustrated in § 4.1.1 and § 4.1.5. A natural choice would be to compute popularity dynamics as close to the actual interaction time as possible. However, experimenting with this can reveal the importance of recent trends in model performance Table 6. Thus, denoting $t-n$ to be $n$ fine time periods ago. We round the long-term horizon according to the change in $t$. Per expectation, we find that $t-6$ and $t-12$ drop performance significantly across all datasets, revealing that for the experimented datasets, capturing trends of the 6 recent weeks is crucial. Interestingly, we find that $t-2$ and $t-3$ perform worse in Epinions, but similarly to $t-1$ in Office and Music. We posit this could be due to dataset-specific cylic patterns where Music and Office may have more short-term recurring trends than Epinions, so not as much information is lost in the previous 2 weeks. This opens up possibilities for other axes which can be used to characterized recommender datasets, independent of size, sparsity, or overall duration.
5.6. Effect of Training Data Available(RQ4)
We investigate the influence of data sparsity on the performance of PrepRec and state-of-the-art sequential recommenders. While varying the training data size, we keep the test data fixed. We remove the oldest interactions from the training set and keep at least 2 training interactions per user to ensure the model is successfully trained.
We can see that the baseline sequential recommenders are sensitive to data sparsity, with significant performance drops when there is less training data. This is expected as the sequential recommenders learn item representations specific to item IDs, and the representation quality is highly dependent on the number of interactions. On the other hand, PrepRec learns item representation through popularity dynamics, showing that popularity dynamics are more robust to data sparsity. We argue that this is crucial for real-world applications where we want to transfer the model to a new domain with limited data. PrepRec is capable of learning from the limited data and outperforms the baseline sequential recommenders.
Figure 2. Recommendation results for varying the number of training interactions. 100% Training percentage corresponds to the result in Table 3. We decrease the training percentage by removing the oldest interactions. PrepRec is robust to data sparsity and outperforms all baseline sequential recommenders when fewer interactions are available.
@@ 5.7. Discussion
PrepRec shows excellent potential as a pre-trained sequential recommendation model, with competitive performance to the best baseline with a fraction of the model size and is capable of zero-shot transfer with minimal performance loss. We argue that PrepRec is particularly useful in the following scenarios: 1) initial sequential model when the data in the domain is sparse; 2) backbone for developing more complex sequential recommenders (i.e., prediction interpolation) 3) online recommendation settings. PrepRec does not explicitly model Item IDs, it can also easily generalize to new items/users with no retraining by just updating the popularity dynamics. We leave this as future work.
We also notice that the training process of state-of-the-art sequential recommenders assumes access to future interactions. For example, Alice has interacted with items $A,B,C$ in the past. The sequential recommenders will use $A,B,C$ to predict the next item $D$. If there’s another person, Bob, whose interaction history is $A,B,C,D$ and all of them happen after Alice’s test interactions $D$, the current sequential recommenders will use Bob’s interaction history $A,B,C,D$ to help Alice predict $D$. This is not realistic in the online recommendation setting. We found out that on average 31% of the training interactions are after the test interactions of the same dataset. More detailed statistics can be found in Appendix E. We argue that this is a major limitation of the current sequential recommenders and should be addressed in the future.
@@ 6. Conclusion
Inspired by the impact of pre-trained models, we propose PrepRec , a pre-trained sequential recommendation model that is capable of cross-domain zero-shot transfer without application-dependent side information. PrepRec learns to represent items based on their popularity dynamics and computes user representations via a popularity dynamics-aware transformer layer. Through extensive experiments, we empirically show that PrepRec can achieve comparable performance to or even outperform state-of-the-art sequential recommenders on regular sequential recommendation tasks. PrepRec also shows excellent generalizability and can zero-shot transfer to a new domain with minimal performance reduction, indicating that popularity dynamics play an important role in sequential recommendation and are crucial for developing generalizable sequential recommenders.
Appendix A Popularity Encoding
We provide details for the item popularity encoding technique introduced in Section 4.1.1. Denote the popularity encoder as $E_{p}$, which takes in a percentile $P(x)\in\mathbb{R}^{+}$. Suppose given the popularity percentile $P(a_{j}^{t})\in\mathbb{R}^{+}$ over a coarse time period $t$, the coarse level popularity vector representation $\mathbf{p}_{j}^{t}\in\mathbb{R}^{k}$ and is computed as follows:
$\mathbf{p}_{j}^{t}$ $=E_{p}(P(a_{j}^{t}))$
$(\mathbf{p}_{j}^{t})_{i}$ =$=\begin{cases}1-\{\frac{P}{10}\},&\text{if }i=\lfloor\frac{P}{10}\rfloor\\ \{\frac{P}{10}\},&\text{if }i=\lfloor\frac{P}{10}\rfloor+1\\ 0,&\text{otherwise}\end{cases}$
[…]
Appendix B Baselines Description
In this section, we provide a brief description of the baselines used in our experiments. We tune the hyper-parameters for each baseline on the validation set and report the best results on the test set.
  • BPR (Rendle et al., 2009) is a learning-to-rank method that learns latent representations for users and items through matrix factorization.
  • MArank (Yu et al., 2019) captures individual and union-level item dependency through multi-order attention and a residual network.
Appendix C Additional Experimental Results
Figure 3. Recommendation results for varying user activity levels. Subsequent groups along the $x$-axis have higher activity levels, i.e., user group 1 corresponds to the least active group of users while group 5 corresponds to the most active ones. We find that PrepRec performs relatively better for users with more interactions
C.1. Performance by User Group
In this experiment, we study the performance of our model on different user activity levels (Figure 3). We divide users into 5 evenly separated groups based on their number of interactions and present the NDCG@10 for each group. We can see from the figure that the gap between baselines and PrepRec decreases as user activity increases for Office and Movie, while PrepRec outperforms baselines for all user groups on Epinions with significant margins. This suggests that in the presence of many interactions, popularity dynamics can capture much of the user’s preferences without using item IDs.
C.2. Parameter Sensitivity
C.2.1. Effect of Window Size
Dataset Music Office Epinions
Metric
R@10
N@10
R@10
N@10
R@10
N@10
36m
0.788
0.529
0.515
0.332
0.751
0.556
24m
0.790
0.553
0.530
0.333
0.748
0.542
0.778
0.553
0.533
0.334
0.757
0.543
6m
0.762
0.546
0.540
0.340
0.761
0.542
3m
0.766
0.545
0.520
0.331
0.764
0.530
2m
0.768
0.528
0.511
0.325
0.745
0.526
1m
0.784
0.543
0.510
0.325
0.747
0.527
Table 7. Recommendation results for varying coarse level popularity dynamics window size $m$ (§ 4.1.1).
We assess the effect of different numbers of window size $m$ used in constructing the coarse level popularity dynamics introduced in § 4.1.1 (Table 7). For all our datasets, we fix $n=4$ for the finer horizon, but experiment with different $m$ for the coarser horizon. The construction method for popularity dynamics and the model’s item embedding sizes remain constant, so we’re effectively trading off incorporating more historical information at the risk of losing nuances in recent trends. Results are presented in Table 7. We find that 6 and 12-month windows are generally best for Office and Epinions, although NDCG@10 improves close to monotonically with window size for Epinions. There isn’t a clear trend for Music, with the 24-month windows performing best, but the 1-month window is competitive as well. This suggests that denser datasets may have more cyclical popularity trends that are captured equally well with shorter and longer windows, while sparser datasets are more sensitive and require more fine-tuning for optimal window size.
Figure 4. Recommendation results for varying the preprocessing discounting factor. $\gamma=0.25$ and $\gamma=0.5$ yield competitive results over the various datasets, while $\gamma=0.75$ is slightly worse.
C.2.2. Effect of Discounting Factor
[…]
C.3. Time and Positional Encoding Ablations
Our model architecture uses a relative time embedding by encoding the relative magnitude time interval between consecutive interactions. Here we assess the importance of this time embedding by comparing the full model, denoted as PrepRec, to a model without it, denoted w/o ${\bm{T}}$. We found that with the addition of relative time encoding ${\bm{T}}$, the performance generally improves. In addition, previous works (Sun et al., 2019; Kang and McAuley, 2018) consider learnable as opposed to fixed positional embeddings. We experiment with learnable positional encodings, denoted Learnable ${\bm{P}}$, for the regular, in-domain setting. We also note that learnable embeddings may be prone to overfitting, and thus won’t generalize well to new domains in zero-shot transfer. We reason that fixed embeddings achieve better performance due to the smaller number of overall parameters our model has (Table 9), resulting in strong test performances for simpler models.
Dataset Music Office
PrepRec N@10
0.553
0.334
w/o ${\bm{T}}$ N@10
0.514
0.334
Learnable ${\bm{P}}$ N@10
0.546
0.330
Table 8. Ablation study for time and positional embeddings. We find that the relative time embedding is important for recommendation performance. On the otherhand, we find that learnable positional embeddings do not perform as well as fixed embeddings.
Appendix D Model Size Analysis
Dataset Office Tool Movie Music Epinions
SasRec
1.331m
3.581m
2.044m
0.542m
1.054m
BERT4Rec
2.687m
7.233m
4.126m
1.094m
2.127m
TiSasRec
1.367m
3.617m
2.127m
0.578m
1.09m
PrepRec
0.045m
0.045m
0.045m
0.045m
0.045m
Table 9. Comparison of model sizes (i.e., number of learnable parameters) over different datasets. PrepRec uses significantly fewer number of parameters than baselines, ranging from 12 to 90x smaller.
In this section, we compare model complexities over different datasets. All the baseline methods learn item embeddings specific to their explicit item IDs, which scale with the number of items in the dataset. This can become a bottleneck for large-scale datasets and prevents zero-shot generalization to new datasets. In contrast, PrepRec learns item representations by modeling their popularity dynamics, which are independent of the number of items and will not scale with the number of items. We compare the model sizes of PrepRec and sequential recommendation baselines in Table 9.
We find that PrepRec uses significantly fewer learnable parameters than sequential recommendation baselines. Indeed, we can see the model size remains constant across datasets of varying sizes, as the number of model parameters does not scale with the dataset, making PrepRec more suitable for large-scale online platforms. Despite its much smaller size, PrepRec is able to achieve competitive performance with baselines.
Appendix E Time Leakage Analysis
Dataset Office Tool Movie Music Epinions
# Future Actions
18.22
18.37
597.56
72.9
27.39
% Future Actions
30.46%
24.39%
45.96%
28.32%
28.70%
Table 10. Time leakage statistics. The first row measures the average number of future interactions of the associated test item included in the training set. The second row measures the average proportion of all of the associated item’s training set interactions that are future interactions.
Sequential recommendation baselines like SasRec, TiSasRec, and BERT4Rec are trained on the users’ sequences without considering the interaction time. As the models train over different user sequences, the item representation of a given item will be updated without any consideration of the exact time at which the interaction occurred. Concretely, since the validation and test sets are relative to each user, a given item $v_{j}$ that interacted with user $u_{i}$ at time $t$ and user $u_{i^{\prime}}$ at time $t^{\prime}$ ($t^{\prime}>t$). The item $v_{j}$ may be part of the training interactions for user $u_{i^{\prime}}$ but the testing interactions for user $u_{i}$. So upon inference time for user $u_{i}$, the item embedding for $v_{j}$ will include information about a future interaction with user $u_{i^{\prime}}$, which would not be available in practice. In contrast, since our method represents items by popularity embeddings up to the time of the interaction, there is no time leakage because there’s no associated item ID, and the training has been conducted only on the universal item representation based on popularity dynamics.
Group
1
2
3
4
5
Consecutive Interactions Time Deviation normal-$\to$
Reg R@10
0.757
0.751
0.781
0.785
0.813
ZS R@10
0.729
0.740
0.771
0.793
0.832
$\Delta$ R@10
-1.5%
-1.2%
+1.0%
Reg N@10
0.518
0.522
0.555
0.567
0.603
ZS N@10
0.4999
0.508
0.546
0.578
0.623
$\Delta$ N@10
-2.6%
-1.6%
+1.9%
First Interaction Time normal-$\to$
Reg R@10
0.795
0.782
0.757
0.762
0.791
ZS R@10
0.780
0.754
0.755
0.771
0.804
$\Delta$ R@10
-1.9%
-0.3%
1.1%
+1.6%
Reg N@10
0.568
0.559
0.541
0.535
0.560
ZS N@10
0.546
0.532
0.531
0.548
0.597
$\Delta$ N@10
-2.0%
2.4%
Last Interaction Time normal-$\to$
Reg R@10
0.656
0.780
0.816
0.750
0.885
ZS R@10
0.614
0.766
0.807
0.795
0.884
$\Delta$ R@10
-1.9%
-1.0%
-0.2%
Reg N@10
0.431
0.548
0.633
0.500
0.652
ZS N@10
0.402
0.529
0.598
0.547
0.679
$\Delta$ N@10
Table 11. Zero-shot Transfer Analysis. We compare the in-domain (trained on target) and zero-shot cross-domain (trained and source and transferred to target) performance of PrepRec. Reg stands for the model performance trained on the target dataset, while ZS stands for the zero-shot cross-domain performance. We use the music as the target dataset and the movie as the source dataset. We divide users into 5 groups based on their interaction time deviation, first interaction time, and last interaction time.
Appendix F Zero-Shot Transfer Analysis
The zero-shot transfer results in Table 4 show that the proposed method can generalize quite well from source to target datasets, achieving competitive performance with baselines and our model trained on the target dataset, despite having no access to the target dataset during training. In this section, we further analyze the performance of zero-shot transfer by comparing its performance to regularly trained PrepRec across different user groups. We consider the Movie dataset as the source and Music dataset as the target, and contrast our regular, in-domain model, denoted Reg with our zero-shot, cross-domain model, denoted ZS. All statistics increase for larger group numbers, as represented by $\to$. $\Delta$ denotes the difference in metrics between ZS and Reg. Statistical significance using the two-sided t-test for independence is indicated by $*{}^{*}$ for $p<0.05$ and ${}^{**}$ for $p<0.01$ in the difference rows.
We find statistically significant results for three different time-based metrics: the deviation in time between consecutive interactions, the time of the first interaction, and the time of the last interaction. Concretely, for each of these metrics, we sort all users by the metric and divide them into 5 groups, with Group 1 being the smallest metric values and Group 5 being the largest, and assess performance. For consecutive interaction deviation, Reg performs better on users with smaller deviations or more frequent interactions, while ZS performs better on users with larger deviations or more spread out interactions. We posit that due to the larger amount of training data, the Movie-trained ZS model captures longer dynamic trends better, but doesn’t cature some shorter-term, idiosyncratic patterns. Also users with earlier first interactions and earlier last interactions also perform better with Reg, while predictions for later first interactions and later last interactions are better with ZS. This could be due to items later in the dataset having more precise popularity data, leading to the pretrained ZS outperforming the in-domain Reg model.

### 10 Apr 2024 v2

@@ 5.4. Regular Sequential Recommendation (RQ1)
We show the performance of PrepRec against the selected baselines. We want to emphasize that the purpose of this paper is not to outperform the state-of-the-art sequential recommenders but to set a basline for further pre-trained sequential recommenders. We include the results here to investigate how well the popularity dynamics can items and sequences in the sequential recommendation. PrepRec achieves competitive performance- within 2% in R@10 and 5% in N@10, - with the state-of-the-art sequential recommenders. On Epinions, PrepRec even outperforms all the sequential recommenders by 7.3%. This is particularly impressive given that PrepRec has significantly fewer model parameters (Table 2).
Since PrepRec explicitly models popularity information and the MostPop demonstrates competitive performance compared to the remaining baselines, we conduct an additional experiment (Table 4) where we sample the unobserved (negative) items based on their popularity (Sun et al., 2019). As shown in Table 4, MostPop’s performance dropped significantly, while PrepRec shows even more competitive performance on some datasets (e.g., Music and Epinions). This suggests even though PrepRec utilizes non-personalized popularity information, it can still learn discriminative item and sequence representations.
@@
PrepRec learns item representations through popularity dynamics, which is conceptually different from learning representations specific to each item ID. Therefore, we propose a simple post-hoc interpolation to investigate if modeling popularity dynamics can boost the performance of state-of-the-art sequential recommenders. We interpolate the scores from PrepRec with the scores from BERT4Rec as follows: $\hat{y}_{intp}(v^{t}_{j}|\mathcal{S}_{u})=\alpha*\hat{y}_{O}(v^{t}_{j}|\mathcal{S}_{u})+(1-\alpha)*\hat{y}_{S}(v_{j}|\mathcal{S}_{u})$, where $\hat{y}_{O}(v^{t}_{j}|\mathcal{S}_{u})$ and $\hat{y}_{S}(v_{j}|\mathcal{S}_{u})$ are the scores from PrepRec (Equation 10) and BERT4Rec, respectively. We set $\alpha=0.5$ for all datasets. After interpolation, there is a significant performance boost by up to 34.9% in N@10. This suggests that our model effectively captures dynamic item properties that are missing with the static item embeddings of SOTA sequential models. Gains are largest in the medium and low-density datasets (Epinions, Amazon), indicating that our model is particularly complimentary to existing methods in sparse datasets where item embeddings are less informative.
5.5. Zero-shot Transfer(RQ2)
We follow the zero-shot inference setting introduced in § 4.3 and report the results in Table 5. We also include the results of the regular sequential recommendation for reference. In the zero-shot setting, after training on source datasets, PrepRec shows minimal performance reduction in the target datasets (i.e., 6% maximum and 2% average reduction in R@10). The best zero-shot transfer results from PrepRec only fall short against the selected sequential recommendation baselines (Table 3) by up to 4% and even outperform (by up to 6.5%) them on the target datasets (e.g., Epinions and Office). We found that our models trained on Movie and tools show the highest generalizability, even outperforming the target-trained models on Music (0.811 vs. 0.782 on R@10). We conjecture that this is because Movie is the largest dataset in terms of the number of interactions. Overall, these results show PrepRec ’s effectiveness in zero-shot transfer without any training on interaction data or side information. In addition, this experiment also demonstrates that the popularity dynamics-based item and sequence representations are generalizable across domains and applications.
@@ 5.6. Ablation Study (RQ3)
PrepRec 0.782 0.573
0.536
0.344 0.795 0.580
w/o Relative Time ${\bm{T}}$ (§ 4.1.3)
0.734
0.514
0.541
0.334
0.782
0.562
w/o Positional ${\bm{P}}$ (§ 4.1.4)
0.765
0.544
0.530
0.332
0.772
0.554
Sinuoisal Popularity Encoding
0.779
0.570
0.529
0.340
0.772
0.561
Table 6. Ablation study of PrepRec ’s different variants.
Here, we assess the importance of different components crucial to PrepRec , i.e., relative time encoding (§ 4.1.3), positional encoding (§ 4.1.4), and popularity encoder $E_{p}$ (§ 4.1.1). We find that removing relative time encoding ${\bm{T}}$ results in the largest performance drop on both the Music and Office datasets. This suggests that the relative time encoding is crucial for effectively capturing the popularity dynamics. Removing positional encoding ${\bm{P}}$ results in a maximum of 2.2% drop in R@10 on the Office dataset, indicating positional encoding is important for capturing sequential information. Finally, changing $E_{p}$ to the non-linear sunusoidal encoding shows worse performance on all datasets, meaning that the linear encoding is more suitable for capturing the popularity dynamics.
5.7. Qualitative Analysis(RQ4)
Figure 2. Recommendation results for different item popularity groups. Item Popularity Group 1 represents the least popular items, and Group 5 represents the most popular items. PrepRec achieves better performance on long-tail items, while the sequential recommenders show stronger performance on popular items.
Here we present a detailed analysis regarding the performance of PrepRec . We separate test items into equally sized groups based on their popularity in the training set. We then compute the average R@10 and N@10 for each group. We present the results in Figure 2. We find that PrepRec achieves better performance on item group with the least interactions, i.e., long-tail items, while the SasRec and BERT4Rec show stronger performance on popular items. Long-tail item recommendation is a particularly challenging task explored by many previous works (Sankar et al., 2021) and requires recommenders able to learn high-quality representations with just a few interactions. This corresponds to our observation that PrepRec is more robust to data sparsity and can learn discriminative item and sequence representations (§ 5.4), showing that long-tail item recommendation can benefit from PrepRec ’s popularity dynamics-based item representations.
5.8. Further Analysis
Dataset Music Office Epinions
Metric
R@10
N@10
R@10
N@10
R@10
N@10
@@ 5.8. Further Analysis
5.8.1. Effect of discounting factor $\gamma$
We examine the effect of different preprocessing weights $\gamma$ used in popularity calculation (§ 4.1.1). In particular, $\gamma=1$ corresponds to the cumulative popularity, or in other words, at a given time period $t$, the overall number of interactions up to period $t$. On the other hand, $\gamma=0$ corresponds to the current popularity, or percentiles are calculated over interactions just in $t$, same as $b_{j}^{t}$ in Equation 1. When $\gamma=0.5$, it can be interpreted as interactions being exponentially weighted by time, with a half-life of 1 time period. We find that $\gamma=0.5$ outperforms the other two settings, with the largest gains of around 12% R@10 and 27% N@10 over cumul-pop in the dense Music dataset, and the largest gains over curr-pop in the sparser Office (5% N@10 and 4% N@10) and Epinions ($15\%$ R@10 and 17%N@10) datasets. We suspect this is due to cumulative measures in denser datasets failing to capture recent trends due to the large historical presence, while current-only measures in sparser datasets convey too little or noisy information and lose the information of long-term trends. curr-pop shows decent performance on the Music dataset, suggesting that Music trends might be more cyclical and thus the current popularity is more informative.
[…]
Metric
R@10
N@10
R@10
N@10
R@10
N@10
PrepRec
0.803
0.591
0.472 0.300 0.489 0.264
SasRec
0.815
0.599
0.437
0.290
0.433
0.245
BERT4Rec
0.816 0.602
0.407
0.249
0.433
0.255
Table 8. Recommendation results for fine-tuning PrepRec. We fine-tune PrepRec and retrain the baselines from scratch on the target dataset.
5.8.2. Fine-tune Capability
While we focus on zero-shot cross-domain and cross-application transfers, we also investigate PrepRec’s fine-tune capability. To ensure target datasets are smaller than the source, we further process the target datasets such that they are no more than 10% of the source datasets’ total interactions. After further processing, we follow the same experimental setup in § 5.3. We fine-tune PrepRec and retrain the baselines from scratch on the target dataset and report the results in Table 8. We find that PrepRec , after fine-tuning, outperforms the selected baselines on Office and Epinions by up to 12.9%. This indicates that PrepRec is capable of learning from the limited data and can be further fine-tuned to achieve better performance, which further expands the use case of PrepRec .
@@ 5.8.2. Fine-tune Capability
Fine:2 days; Coarse:10 days $*{}^{*}$
0.782 0.573
0.536
0.344 0.795 0.580
Fine:4 days; Coarse:15 days
@@
Table 9. Recommendation results for varying time horizons. Fine and coarse time horizons are used for short-term and long-term popularity dynamics respectively (§ 4.1.1).
5.8.3. Effect of Different Time Horizons
We study the effect of different time horizons to PrepRec . We found that in general, long-term horizons of 30 days and short-term horizons of 7 days perform worse than the other settings. This is likely because the long-term horizon might lead to the lack of resolutions in popularity statistics. We also find that depending on the dataset, the effect of different time horizons also varies. For example, both Music and Epinions show larger performance decrease from short to long-term horizons than Office. This could be because Music and Epinions are more sensitive to recent trends than Office, or their data are denser in terms of time granularity.
Dataset Music Office Epinions
Metric
R@10
N@10
R@10
N@10
R@10
N@10
0.782 0.573 0.536 0.344 0.795 0.580
@@ 5.8.4. Effect of Prediction Time
In PrepRec, we use the popularity trends up to some time as the context to construct universal item representations. To prevent information leakage, we can only use the popularity dynamics computed prior to the time of the interaction, as illustrated in § 4.1.2 and § 4.1.6. Intuitively, we should compute popularity dynamics as close to the actual interaction time as possible. However, experimenting with this can reveal the importance of recent trends in model performance ( Table 10). Thus, denoting $t-n$ to be $n$ fine time periods ago. We round the long-term horizon according to the change in $t$. Per expectation, we find that $t-6$ and $t-12$ drop performance significantly across all datasets while $t-2$ performs relatively close to $t-1$, revealing that for the experimented datasets, capturing the latest popularity changes is crucial.
@@ 5.9. Discussion
PrepRec shows excellent potential as a pre-trained sequential recommendation model, with competitive performance to the best baseline with a fraction of the model size, and is capable of zero-shot transfer across domains and applications. We argue that PrepRec is particularly useful in the following scenarios: 1) initial sequential model when the data in the domain is sparse; 2) backbone for developing more complex sequential recommenders (i.e., prediction interpolation) 3) online recommendation settings. PrepRec does not explicitly model Item IDs, it can also easily generalize to new items/users with no retraining by just updating the popularity dynamics. We leave this as future work.
@@ 6. Conclusion
Inspired by the impact of pre-trained models, we propose PrepRec , a pre-trained sequential recommendation model that is capable of cross-domain and cross-application zero-shot transfer. We assume no access to auxiliary information (e.g., textual data), and we set a baseline for pre-trained sequential recommenders under the most challenging setting. PrepRec learns to represent items based on their popularity dynamics and computes sequence representations via a popularity dynamics-aware transformer. Through extensive experiments, we empirically show that PrepRec can achieve comparable performance to or even outperform state-of-the-art sequential recommenders on regular sequential recommendation tasks. PrepRec also shows excellent generalizability and can zero-shot transfer to a new domain with minimal performance reduction, indicating that popularity dynamics play an important role in sequential recommendation and are crucial for developing generalizable sequential recommenders.